OpenStudy (anonymous):

p(-4,2), y=-5f(4x)-8 find the corresponding point on the graph to the function.

OpenStudy (anonymous):

So I thought x would be 16, and y -18? but it was wrong anyone walk be through this?

OpenStudy (anonymous):

@wio

OpenStudy (tkhunny):

We know f(-4) = 2. This is ALL we know. We know NOTHING of this function except its value at -4. 4x is the argument in the new expression. If we pick x = -1, we have y = 5f(4(-1)) - 8 = 5f(-4) - 8 = 5(2) - 8 = 10 - 8 = 2 ...and we have (x = -1 , y = 2) Note: We do NOT have f(-1) = 2 -- Not even a little.

OpenStudy (anonymous):

Still lost.

OpenStudy (tkhunny):

Lost by what? Can you substitute into function notation? I did that. So can you. Can you replace an expression by its value? I did that. So can you. Can you manipulate a small collection of numbers and operations? I did that. So can you. So, what is it that causes you to be lost?

OpenStudy (anonymous):

Ok so I get f(-4)=2 I am lost on why we picked x=-1 I though we take the x (so -4) so substitute it in (4(-4)) and that give you the first point? -16 and for the second point -5 ( 2 ) =-10 now subtract the -8 giving you -18 so the points would be (-16,-18)

OpenStudy (tkhunny):

You do not understand the importance of my first little paragraph, repeated here for convenience: We know f(-4) = 2. This is ALL we know. We know NOTHING of this function except its value at -4. We do NOT know f(-16). The ONLY thing we know is f(-4). We do NOT know f(4(-4)). We just don't. The ONLY thing we know is f(-4). The ONLY way to get f(4) out of f(4x) is to pick x = -1.